OpenStudy (anonymous):

Is (6,-6) a solution of 2y-4x=-12

Mehek (mehek14):

plug in the values 2x6 - 4x-6 = -12 12 + 24 = 36 no it is not a solution it would be if the ordered pairs were (6,6)

OpenStudy (anonymous):

Thank you sooooo much!!! You are awesome

Mehek (mehek14):

wait hold on 2x-6 - 4x6 = -12 -12 - 24 = -36 either way no it isn't

Mehek (mehek14):

i switched the numbers in the beginning

OpenStudy (anonymous):

Your answer is correct

OpenStudy (anonymous):

Is (0,0) a solution to y=-2x?